# TALE: A Tool-Augmented Framework for Reference-Free Evaluation of Large Language Models

### v1

@@ 1 Introduction
Consider the question: “What is the tallest building in the world?” The correct answer has changed over time—from the “Willis Tower” to “Taipei 101”, and more recently, to the “Burj Khalifa.” and could be expressed in various valid ways: “Burj Khalifa,” “Khalifa Tower,” “Burj Dubai,” or with its precise height. Traditional reference-based evaluation metrics, including Exact Match (EM), F1, BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004), rely on fixed ground-truth references and thus may incorrectly penalize such legitimate variations (Kamalloo et al., 2023; Wang et al., 2023). These methods, though efficient, struggle to reflect semantic equivalence, are limited in adaptability to evolving facts. On the other hand, human evaluation is generally reliable, but it becomes overly expensive and labor-intensive when applied at scale (Chiang & Lee, 2023; Mañas et al., 2024; Zhu et al., 2023).
The challenge of evaluating LLMs grows as they act as autonomous agents, handling tasks like web browsing, planning, and information synthesis (Xi et al., 2023; Fu et al., 2023). In such dynamic settings, outputs are often unpredictable, context-dependent, and non-deterministic, making it impractical to pre-annotate reference answers for every possible interaction (Li et al., 2024). As a result, static, reference-driven evaluation protocols are fundamentally misaligned with the needs of modern LLM applications operating in open-ended, real-world environments.
@@
Our key contributions include:
  • We introduce TALE, a novel framework that augments LLM-based evaluation with iterative, tool-assisted evidence gathering to overcome the limitations of both reference-based metrics and ungrounded LLM self-judgment.
  • We implement and evaluate a multi-step reasoning process, enabling LLM judges to actively verify factual claims rather than relying solely on internal knowledge
  • We demonstrate through extensive experiments that TALE is aligned with reference-based metrics such as F1 and achieves substantial to perfect agreements with human evaluators.
  • We provide ablation studies that quantify the impact of iterative refinement and query generation capabilities on evaluation performance. The results offer insights into the key components that drive TALE’s effectiveness.
@@ 2.3 Tool-Augmented Evaluation
An emerging category attempts to overcome the limitations of closed-world LLM evaluators by incorporating external tools. WebGPT (Nakano et al., 2021) demonstrated how web-browsing capabilities improve factual correctness, while SelfCheckGPT (Manakul et al., 2023) used multiple independent generations to perform self-consistency checks. More closely related to our work, FActScore (Min et al., 2023) decomposes generated text into atomic facts and verifies each against retrieved external knowledge.
While these approaches leverage external information, they differ fundamentally from TALE in several aspects. First, most focus on evaluation through fact extraction and verification rather than holistic response assessment. Second, they typically perform single-pass evidence retrieval without the iterative refinement that characterizes TALE. Finally, they often lack explicit reasoning about what information to retrieve, instead relying on keyword-based or embedding similarity retrievals.
Our TALE framework addresses these gaps by implementing an Agent-as-a-judge paradigm (Zhuge et al., 2024) with iterative, adaptive evidence gathering, explicit reflection, and comprehensive reasoning, enabling more robust evaluation in scenarios where static reference answers are unavailable or insufficient.
@@ 3 Methodology
Figure 1: Given the question, “What is the tallest building in the world?” and the candidate answer, “The tallest building in the world is the Burj Khalifa in Dubai.”, the process begins with an initial query generation based solely on the question to extract relevant information. This query is used for evidence gathering through web searches across multiple sources. Once the evidence is collected, it is processed by the summary module, which condenses it and extracts key insights. The summarized evidence is then passed to the reflector module, which assesses the sufficiency and relevance of the evidence. In case of insufficient or conflicting evidence, it triggers a query refinement step to improve the evidence through further iterations. After N iterations, the system enters the judgment phase where the judge synthesizes all available evidence and provides a final decision with rationale.
@@ 3.2 LLM-as-a-Judge
In our framework, an LLM is employed as a judge, denoted by $J$, to assess the correctness of $\hat{y}$ based on external evidence. Instead of relying on a fixed reference, the judge acts as an autonomous agent that leverages the aggregated evidence $E$ and reflection $R$, obtained through an iterative process, along with the input question $x$ and the candidate answer $\hat{y}$.
Formally, the judge produces a binary verdict $v\in\{0,1\}$ (i.e., True or False) and a brief natural language rationale $r_{J}$, such that:
$(v,r_{J})=J(x,\hat{y},E,R).$
Here, $J$ is a mapping defined as
$J:\mathcal{X}\times\mathcal{Y}\times\mathcal{E}\rightarrow\{0,1\}\times\mathbb{R},$
where $v=1$ indicates that the candidate answer $\hat{y}$ is deemed correct based on the evidence $E$, and $v=0$ otherwise. The judge gives precedence to evidence gathered from external sources but may fall back on their pre-trained knowledge when the evidence is incomplete or inconclusive.
@@ 3.3 Tool-Augmented Evaluation Process
In the first iteration ($i=1$), the query $q_{1}$ is constructed solely from the input question $x$: $q_{1}=Q(x)$. The goal is to retrieve background information relevant to the question domain, without assuming the correctness of any particular answer. For example, given the question “What is the capital of Switzerland?”, the model may generate a search query like “Capital of Switzerland” or “Switzerland official capital city”. This ensures that the evaluation process begins with a neutral, answer-agnostic evidence-gathering step.
In subsequent iterations ($i>1$), queries are refined using accumulated evidence and reflections stored in the short-term memory: $q_{i}=Q^{\prime}(\mathcal{M})$. The purpose is to resolve remaining uncertainty or retrieve more targeted information needed to verify or refute the candidate answer $\hat{y}$. For instance, if the candidate’s answer claims “Bern is the capital of Switzerland” but initial evidence suggests there might be a distinction between political and economic capitals, a refined query might be “Is Bern the official capital of Switzerland or is there another capital”.
@@
The query $q_{i}$ is submitted to a web search via the Serper API, [^1] https://serper.dev/ to return real-time results. For each query, we retrieve the top $k=3$ search snippets or URLs, which serve as raw external evidence for the subsequent steps. The search results typically include titles, snippets, and source URLs, providing diverse perspectives from multiple information sources. This module enables TALE to access up-to-date and diverse information sources beyond the model’s pre-trained knowledge.
@@
The retrieved search results—comprising multiple web snippets or document passages—are passed to a summarization module that condenses them into a focused evidence segment: $E_{i}=\Sigma(S(q_{i}))$. This module extracts salient factual content while filtering out redundant or irrelevant information. The resulting summary $E_{i}$ serves as a concise, interpretable knowledge unit that informs reflection. For example, a summary might state: “Multiple sources confirm that Bern is the de facto capital and seat of the federal government of Switzerland, though Switzerland does not have an official capital city designated by its constitution.” In TALE, this step is critical for managing context length and ensuring that only high-signal evidence is retained throughout the iterative reasoning process.
@@
The summarized evidence $E_{i}$ is evaluated in relation to the input question $x$ and the candidate answer $\hat{y}$ to assess its relevance, sufficiency, and factual alignment. This step yields a reflection signal: $R_{i}=R(x,\hat{y},E_{i})$, which captures whether the current evidence supports, contradicts, or is inconclusive with respect to $\hat{y}$.
The reflection module also identifies missing information or ambiguities that can guide subsequent query refinement. By explicitly reasoning over the current evidence, reflection enables TALE to iteratively improve its evaluation behavior over multiple steps.
@@
After each iteration, the tuple $(q_{i},E_{i},R_{i})$—consisting of the generated query, corresponding evidence, and reflection—is appended to the short-term memory buffer:
$\mathcal{M}\leftarrow\mathcal{M}\cup\{(q_{i},E_{i},R_{i})\}$
This memory $\mathcal{M}$ serves as an evolving trace of the evaluation process, accumulating context that informs subsequent query refinement and reasoning. The memory is append-only and scoped to a single evaluation episode, ensuring that each step builds upon the full history of prior actions, observations, and reflections.
@@
After $N$ iterations, the judge module synthesizes all gathered evidence to assess the correctness of the candidate answer. Specifically, it includes the aggregated evidence summary and reflections. Based on such prior components, $J$ produces a binary verdict $v\in\{0,1\}$ indicating whether $\hat{y}$ is factually correct, along with a natural language rationale $r_{J}$:
$(v,r_{J})=J(x,\hat{y},E_{\text{total}},R_{\text{total}}).$
The overall procedure is summarized in Algorithm 1.
@@ 4 Experimental Setup
This section details our experimental setup for evaluating TALE across multiple dimensions.
@@ 4.1 Models
In our experiments, we utilize: Gemini-1.5-pro (Team, 2024), GPT-3.5-turbo (Brown et al., 2020), and GPT-4o-mini (Team, 2023), both as candidates ($C$) and as judges ($J$) within the TALE framework, allowing us to assess both their ability to generate accurate responses and to evaluate responses from other models. Additionally, we also explore the potential of open-source LLMs as judges, including Mistral 7B [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). All experiments are conducted with a temperature of 00 to maximize determinism and reliability, as prior work demonstrates that higher temperatures degrade the performance of LLM-based evaluators (Hada et al., 2024). For brevity, we refer to these models as Mistral, Llama, Gemini, GPT-3.5, and GPT-4o throughout our analysis.
@@ 4.2 Datasets
We evaluate TALE on widely used free-form question-answering datasets that represent different question types, knowledge domains, and complexity levels. These includes AmbigQA (Min et al., 2020), HotpotQA (Yang et al., 2018), TriviaQA (Joshi et al., 2017), and Natural Questions (NQ-Open) (Kwiatkowski et al., 2019). Free-form question-answering underpins a broad range of practical applications, where maintaining accuracy and ensuring truthfulness are paramount (Gou et al., 2024). We also used FreshQA (Vu et al., 2023) to see TALE’s ability in detecting outdated knowledge. Due to computational constraints, we randomly sampled 300 instances from each dataset to ensure balanced representation across question types and difficulty levels. Each dataset provides reference answers that serve as ground truth for our reference-based baseline metrics (details in Appendix A.1).
@@ 4.3 Prompts
Our prompting strategy uses templates for both response generation and evaluation. For candidate models, we use few-shot Chain-of-Thought (CoT) prompts with 6 examples per dataset to elicit detailed, reasoning-based responses.
For the TALE framework, we design module-specific prompts that combine role instructions with step-by-step reasoning guidance. Each module (query generation, evidence summarization, reflection, refinement, and judgment) uses a one-shot prompt template that demonstrates the expected input-output behavior. These prompts are crafted to encourage explicit reasoning and evidence-focused analysis without assuming the correctness of any particular answer. Complete prompt templates and examples are provided in Appendix A.2.
@@ 4.4 Baselines
We compare TALE against several established evaluation approaches, including reference-based metrics and a reference-free judge. For the reference-based metrics, we implement two widely-used automatic metrics, Exact Match (EM) and F1. Additionally, we adopt a judge without tool-augmentation (G-Eval) as a baseline, following the approach from Liu et al. (2023), where the judge LLM evaluates candidate answers based solely on the question-answer pair without access to external tools or reference answers. In this setting, the judge relies entirely on its pre-trained knowledge to determine factual correctness. By maintaining the same judge model while removing the evidence retrieval mechanism, this baseline isolates the impact of tool augmentation (details in Appendix A.3).
@@ 4.5 Evaluation Metrics
To assess TALE’s performance, we used accuracy as the proportion of instances where the judge’s binary verdict (correct/incorrect) aligns with the ground truth derived from reference answers. For AmbigQA and HotpotQA that include human annotations, we calculate Cohen’s Kappa ($\kappa$) and Macro-F1 scores to evaluate TALE’s alignment with human majority votes, with Kappa accounting for agreement beyond chance and Macro-F1 addressing class balance. Additionally, we conduct ablation studies to quantify the impact of removing or modifying specific TALE components, using changes in agreement with human judgments as the primary measure. For component ablation studies, we measure the change in agreement with human judgments when removing or modifying specific TALE modules.
@@ 5 Results
This section briefly presents the results obtained from our experiments. We included additional results and analysis in the Appendix B.
@@
Table 1: Raw performance of candidate LLMs obtained through different evaluators. Evaluation approaches include: (1) Reference-based metrics (EM, F1), (2) Judge w/o Tool (accuracy), and (3) TALE (accuracy).
TALE works better with more powerful LLMs.
@@
To evaluate TALE’s ability to detect outdated knowledge, we included FreshQA (Vu et al., 2023), Based on a random sample of 300 questions, we found that TALE consistently identified outdated information in candidate model responses. For instance, when asked “Where is EMNLP this year?”, candidate models often provided outdated responses based on their training data. TALE retrieved current information indicating the correct location, “Suzhou, China.” (see Appendix B.3).
@@
5.1 Error analysis
To better understand the limitations of TALE in evaluating candidate responses, we conducted a manual error analysis. We randomly sampled 100 evaluation cases from the AmbigQA and HotpotQA datasets, focusing on instances where TALE disagreed with human annotators. We categorized the errors into the following categories: 1) Contextual misunderstanding: TALE generates inaccurate or incomplete queries when it misinterprets the intent of the candidate’s question. This is particularly evident in AmbigQA, where questions are often intentionally ambiguous or lack sufficient context, leading to the retrieval of irrelevant or contradictory evidence. 2) Incomplete evidence: TALE fails when the retrieved evidence is insufficient or lacks relevant information, specifically for recent events with limited online coverage. 3) Reasoning error: Despite accurate evidence, the judge model misinterprets the information or applies flawed reasoning. 4) Hallucination: In cases where evidence is ambiguous or inconclusive, TALE relies on its pre-trained knowledge, resulting in hallucinated rationales. 5) Conflicting evidence: In some cases, TALE encounters conflicting evidence across multiple search iterations. While the framework is designed to iteratively refine its understanding and gather reliable information, judges sometimes over-rely on earlier sources or fail to appropriately weigh the credibility of conflicting information (details in Appendix B.5).

### v2

@@ 1 Introduction
Traditional evaluation methods, including human evaluation and supervised signal-based metrics such as Exact Match (EM), F1, BLEU (Papineni et al., 2002), or ROUGE (Lin, 2004), rely on comparisons to static ground-truth references. While convenient and efficient, these methods fall short of capturing the diversity of valid outputs and are often infeasible to scale due to the high cost of human annotations (Chiang & Lee, 2023; Mañas et al., 2024; Zhu et al., 2023). The challenge of evaluating LLMs output grows as they act as autonomous agents, handling tasks like web browsing, planning, and information synthesis (Xi et al., 2023; Fu et al., 2023). In such dynamic settings, outputs are often unpredictable, context-dependent, and non-deterministic, making it impractical to pre-annotate reference answers for every possible interaction (Li et al., 2024). As a result, static, reference-driven evaluation protocols are fundamentally misaligned with the needs of modern LLM applications operating in open-ended, real-world environments.
@@
Our key contributions include: (1) We introduce TALE, a novel framework that augments LLM-based evaluation with iterative, tool-assisted evidence gathering to overcome the limitations of both reference-based metrics and ungrounded LLM self-judgment. (2) We implement and evaluate a multi-step reasoning process, enabling LLM judges to actively verify factual claims rather than relying solely on internal knowledge. (3) We demonstrate through extensive experiments that TALE is aligned with reference-based metrics such as F1 and achieves substantial to perfect agreements with human evaluators.
@@ 2.3 Tool-Augmented Evaluation
An emerging category attempts to overcome the limitations of closed-world LLM evaluators by incorporating external tools. WebGPT (Nakano et al., 2021) demonstrated how web-browsing capabilities improve factual correctness, while SelfCheckGPT (Manakul et al., 2023) used multiple independent generations to perform self-consistency checks. More closely related to our work, FActScore (Min et al., 2023) decomposes generated text into atomic facts and verifies each against retrieved external knowledge. Concurrently, agent-as-a-judge (Zhuge et al., 2024) introduces a tool-augmented judge-agent that walks through an agent’s entire execution trace and labels hundreds of explicit requirements. It reaches near-human accuracy while reducing evaluation cost by roughly two orders of magnitude.
While these approaches leverage external information, they differ fundamentally from TALE in several aspects. Most importantly, previous LLM-as-a-Judge methods output subjective plausibility or preference scores, whereas TALE delivers objective true/false decisions, each accompanied by a rationale grounded in an automatically assembled, evidence-backed checklist. Beyond this, earlier approaches score isolated facts rather than the full answer, issue a one-shot evidence query instead of TALE’s iterative retrieve–refine loop, and treat retrieval as a keyword-matching black box rather than a planned, reasoning-driven step.
@@ 3 Methodology
Figure 1: Given the question “What is the tallest building in the world?” and candidate answer “The tallest building in the world is the Burj Khalifa in Dubai,”, TALE begins with initial query generation from the question alone. The query triggers web searches across multiple sources, followed by evidence summarization to extract key insights. The reflector module assesses evidence sufficiency and relevance, triggering query refinement if needed. After N iterations, the judgment phase synthesizes all evidence to provide a final decision with rationale.
@@ 3.2 LLM-as-a-Judge
In our framework, an LLM is employed as-a-judge, denoted by $J$, to assess the correctness of $\hat{y}$ based on external evidence.
@@ 3.3 Tool-Augmented Evaluation Process
In the first iteration ($i=1$), the query $q_{1}$ is constructed solely from the input question $x$: $q_{1}=Q(x)$. The goal is to retrieve background information relevant to the question domain, without assuming the correctness of any particular answer. In subsequent iterations ($i>1$), queries are refined using accumulated evidence and reflections stored in the short-term memory: $q_{i}=Q^{\prime}(\mathcal{M})$. The purpose of query refinements is to resolve remaining uncertainty or retrieve more targeted information needed to verify or refute the candidate answer $\hat{y}$.
@@
The query $q_{i}$ is submitted to a web search via the Serper API, [^1] https://serper.dev/ to return real-time results. For each query, we retrieve the top $k=3$ search snippets or URLs, which serve as raw external evidence for the subsequent steps. The search results typically include titles, snippets, and source URLs, providing diverse perspectives from multiple information sources.
@@
The retrieved search results—comprising multiple web snippets or document passages—are passed to a summarization module that condenses them into a focused evidence segment: $E_{i}=\Sigma(S(q_{i}))$. This module extracts salient factual content while filtering out redundant or irrelevant information. The resulting summary $E_{i}$ serves as a concise, interpretable knowledge unit that informs reflection.
@@
The summarized evidence $E_{i}$ is evaluated in relation to the input question $x$ and the candidate answer $\hat{y}$ to assess its relevance, sufficiency, and factual alignment. This step yields a reflection signal: $R_{i}=R(x,\hat{y},E_{i})$, which captures whether the current evidence supports, contradicts, or is inconclusive with respect to $\hat{y}$. The reflection module also identifies missing information or ambiguities that can guide subsequent query refinement. By explicitly reasoning over the current evidence, reflection enables TALE to iteratively improve its evaluation behavior over multiple steps.
@@
After each iteration, the tuple $(q_{i},E_{i},R_{i})$—consisting of the generated query, corresponding evidence, and reflection—is appended to the short-term memory buffer: $\mathcal{M}\leftarrow\mathcal{M}\cup\{(q_{i},E_{i},R_{i})\}$. This memory $\mathcal{M}$ serves as an evolving trace of the evaluation process, accumulating context that informs subsequent query refinement and reasoning. The memory is append-only and scoped to a single evaluation episode, ensuring that each step builds upon the full history of prior actions, observations, and reflections.
@@
After $N$ iterations, the judge module synthesizes all gathered evidence to assess the correctness of the candidate answer. Specifically, it includes the aggregated evidence summary and reflections. Based on such prior components, $J$ produces a binary verdict $v\in\{0,1\}$ indicating whether $\hat{y}$ is factually correct, along with a natural language rationale $r_{J}$: $(v,r_{J})=J(x,\hat{y},E_{\text{total}},R_{\text{total}})$. The overall procedure is summarized in Algorithm 1.
@@ 4.1 Models
In our experiments, we utilize: Gemini-1.5-pro (Team, 2024), GPT-3.5-turbo (Brown et al., 2020), and GPT-4o-mini (Team, 2023), both as candidates ($C$) and as judges ($J$) within the TALE framework, allowing us to assess both their ability to generate accurate responses and to evaluate responses from other models. Additionally, we also explore the potential of open-source LLMs as judges, including Mistral 7B [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). All experiments are conducted with a temperature of 00 to maximize determinism and reliability, as prior work demonstrates that higher temperatures degrade the performance of LLM-based evaluators (Hada et al., 2024). For brevity, we refer to these models as Mistral, Gemini, GPT-3.5, and GPT-4o throughout our analysis.
@@ 4.2 Datasets
We evaluate TALE on widely used free-form question-answering datasets that represent different question types, knowledge domains, and complexity levels. These includes AmbigQA (Min et al., 2020), HotpotQA (Yang et al., 2018), TriviaQA (Joshi et al., 2017), and Natural Questions (NQ-Open) (Kwiatkowski et al., 2019). Free-form question-answering underpins a broad range of practical applications, where maintaining accuracy and ensuring truthfulness are paramount (Gou et al., 2024). We also used FreshQA (Vu et al., 2023) to see TALE’s ability in detecting outdated knowledge. Due to computational constraints, we randomly sample 300 instances from each dataset, ensuring balanced representation across question types and difficulty levels. Each dataset provides reference answers that serve as ground truth for our reference-based baseline metrics (see Appendix A.1).
@@ 4.3 Prompts
Our prompting strategy uses templates for both response generation and evaluation. For candidate models, we use few-shot Chain-of-Thought (CoT) prompts with 6 examples per dataset to elicit detailed, reasoning-based responses. For the TALE framework, we design module-specific prompts that combine role instructions with step-by-step reasoning guidance. Each module (query generation, evidence summarization, reflection, and judgment) uses a one-shot prompt template that demonstrates the expected input-output behavior. These prompts are crafted to encourage explicit reasoning and evidence-focused analysis without assuming the correctness of any particular answer. See Appendix A.2 for examples.
@@ 4.4 Baselines
We compare TALE against several established evaluation approaches, including reference-based metrics and a reference-free judge. For the reference-based metrics, we implement two widely-used automatic metrics, Exact Match (EM) and F1. Additionally, we adopt a judge without tool-augmentation (G-Eval) as a baseline, following the approach from Liu et al. (2023). In this setting, the judge relies entirely on its pre-trained knowledge to determine factual correctness. By maintaining the same judge model while removing the evidence retrieval mechanism, this baseline effectively isolates the impact of tool augmentation in TALE (details in Appendix A.3).
@@ 4.5 Evaluation Metrics
To assess TALE’s performance, we used accuracy as the proportion of instances where the judge’s binary verdict (correct/incorrect) aligns with the ground truth derived from reference answers. For AmbigQA and HotpotQA that include human annotations, we calculate Cohen’s Kappa ($\kappa$) and Macro-F1 scores to evaluate TALE’s alignment with human majority votes. Additionally, we conduct ablation performance studies to quantify the impact of removing or modifying specific TALE components, using changes in agreement with human judgments as the primary measure.
@@ 5 Results
This section briefly presents the results obtained from our experiments. We included additional results and analysis in the Appendix B and further ablations in the Appendix C.
@@
TALE works better with more powerful LLMs.
@@
Table 1: Raw performance of candidate LLMs obtained through different evaluators.
@@
To evaluate TALE’s ability to detect outdated knowledge, we included FreshQA (Vu et al., 2023). Based on a random sample of 300 questions, we found that TALE consistently identified outdated information in candidate model responses. For instance, when asked “Where is EMNLP this year?”, candidate models often provided outdated responses based on their training data. TALE retrieved current information indicating the correct location, “Suzhou, China.” (see Appendix B.3).
@@
5.1 Error analysis
To better understand the limitations of TALE in evaluating candidate responses, we conducted a manual error analysis. We randomly sampled 100 evaluation cases from the AmbigQA and HotpotQA datasets, focusing on instances where TALE disagreed with human annotators. We categorized the errors into the following categories: 1) Contextual misunderstanding: TALE generates inaccurate or incomplete queries when it misinterprets the intent of the candidate’s question. This is particularly evident in AmbigQA, where questions are often intentionally ambiguous or lack sufficient context, leading to the retrieval of irrelevant or contradictory evidence. 2) Incomplete evidence: TALE fails when the retrieved evidence is insufficient or lacks relevant information, specifically for recent events with limited online coverage. 3) Reasoning error: Despite accurate evidence, the judge model misinterprets the information or applies flawed reasoning. 4) Hallucination: In cases where evidence is ambiguous or inconclusive, TALE relies on its pre-trained knowledge, resulting in hallucinated rationales. 5) Conflicting evidence: In some cases, TALE encounters conflicting evidence across multiple search iterations. While the framework is designed to iteratively refine its understanding and gather reliable information, judges sometimes over-rely on earlier sources or fail to appropriately weigh the credibility of conflicting information (details in Appendix C.3).